Shift report 20000801 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Aug 01 2000 - 23:45:20 EDT

  • Next message: http account: "Shift report 20000802 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Aug  1 23:45:20 2000, Shift: 16:00-24:00
    Supervisor: GL, with: +oh
    Log book#3 pages 124 - 128
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam. Closed ring @ 1800, since then "try to restore
       blue ring".
    2. GL and OH learned a lot about restoring HV (see Konstan-
       tin's emails to night) and magnets (ask the watch!).
       1 small frog evicted @ 2021 hours
    3. a) T2 still at an uncertain status
       b) T4 taken out, repaired and back in
       c) FFS 1440 mainframe "dead as a duck"
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Aug 01 2000 - 23:48:57 EDT